OpenStudy (anonymous):

Find the exact answer. log8(7x+1) = -2

OpenStudy (imstuck):

That 8 is the base of the log, right?

OpenStudy (anonymous):

yes

OpenStudy (imstuck):

ok, watch and learn, my dear...

OpenStudy (anonymous):

alrighty

OpenStudy (anonymous):

log8 (x²-7x)=1 x²-7x = 8 x²-7x-8 = 0 x²-8x+x-8 = 0 x(x-8)+1(x-8) = 0 (x+1)(x-8) = 0 x = {-1,8}

OpenStudy (imstuck):